A ceiling fan has a diameter ( of the circle through the outer edges of the three blades ) of 120 cm & rpm 1500 at full speed. Consider a particle of mass 1 g sticking at the outer end of a blade. How much force does it experience when the fan runs at full speed ? Who exerts this force on the particle ? How much force does the particle exert on the blade along its surface ?


Grade:upto college level

1 Answers

Suparshwa Patil
7 years ago
Diameter=120 cmTherefore, radius=60cm=0.6mRpm(n)=1500 Therefore revolution per second=1500/60=25Omega(w)=2pien=2*3.14*25=157 rads/secForce=mw^2r=0.001*157*157*0.6=14.8N